# Camera-Pose Robust Crater Detection from Chang’e 5

### v1

@@ Abstract
As space missions aim to explore increasingly hazardous terrain, accurate and timely position estimates are required to ensure safe navigation. Vision-based navigation achieves this goal through correlating impact craters visible through onboard imagery with a known database to estimate a craft’s pose. However, existing literature has not sufficiently evaluated crater-detection algorithm (CDA) performance from imagery containing off-nadir view angles. In this work, we evaluate the performance of Mask R-CNN for crater detection, comparing models pretrained on simulated data containing off-nadir view angles and to pretraining on real-lunar images. We demonstrate pretraining on real-lunar images is superior despite the lack of images containing off-nadir view angles, achieving detection performance of 63.1% $F_{1}$-score and ellipse-regression performance of 0.701 intersection over union. This work provides the first quantitative analysis of performance of CDAs on images containing off-nadir view angles. Towards the development of increasingly robust CDAs, we additionally provide the first annotated CDA dataset with off-nadir view angles from the Chang’e 5 Landing Camera, available here.
@@ 2 Related Works
There exists significant literature describing the creation and evaluation of CDAs under different operating conditions [tewari2023, Woicke2018-lg, DeLatte2019-ye, Salamuniccar2008-oy]. Typically, CDAs focus on detecting lunar or Mars craters, towards the goal of increasing completeness of crater catalogues of these planetary bodies [Benedix2020, Lee2021, AliDib2020]. As such, many CDAs use DEMs or digital-terrain maps (DTMs) as input, as this data is free from variation due to lighting or view angle [tewari2023]. However, for optical navigation, crater detection from a DEM or DTM is constraining as these maps require many measurements to be accurately defined [LRO-Dataset]. Hence, for optical navigation, it is preferable to detect craters from optical imagery.

### v2

@@ Abstract
As space missions aim to explore increasingly hazardous terrain, accurate and timely position estimates are required to ensure safe navigation. Vision-based navigation achieves this goal through correlating impact craters visible through onboard imagery with a known database to estimate a craft’s pose. However, existing literature has not sufficiently evaluated crater-detection algorithm (CDA) performance from imagery containing off-nadir view angles. In this work, we evaluate the performance of Mask R-CNN for crater detection, comparing models pretrained on simulated data containing off-nadir view angles and to pretraining on real-lunar images. We demonstrate pretraining on real-lunar images is superior despite the lack of images containing off-nadir view angles, achieving detection performance of 63.1% $F_{1}$-score and ellipse-regression performance of 0.701 intersection over union. This work provides the first quantitative analysis of performance of CDAs on images containing off-nadir view angles. Towards the development of increasingly robust CDAs, we additionally provide the first annotated CDA dataset with off-nadir view angles from the Chang’e 5 Landing Camera, available here: https://zenodo.org/doi/10.5281/zenodo.11326449.
@@ 2 Related Works
There exists significant literature describing the creation and evaluation of CDAs under different operating conditions [tewari2023, Woicke2018-lg, DeLatte2019-ye, Salamuniccar2008-oy]. Typically, these algorithms aim to increase the completeness of crater catalogues on planetary bodies such as the Moon or Mars by automating detection [Benedix2020, Lee2021, AliDib2020]. As such, many CDAs use DEMs as input, as this data is free from variation due to lighting or view angle [tewari2023]. However, for optical navigation, crater detection from a DEM is constraining as these maps require many measurements to be accurately defined [LRO-Dataset]. Hence, for optical navigation, it is preferable to detect craters from optical imagery.